# Diffusion Models Meet Contextual Bandits

### v2

@@ 4.1 Discussion
Large action space aspect and regret independent of $K$? dTS’s regret bound scales with $K\sigma_{1}^{2}$ instead of $K\sum_{\ell}\sigma_{\ell}^{2}$, which is particularly beneficial when $\sigma_{1}$ is small, as often seen in diffusion models. Both our regret bound and experiments demonstrate that dTS outperforms LinTS more significantly as the action space grows. Previous studies [25, 67, 70] proposed bandit algorithms whose regret do not scale with $K$, but our setting is fundamentally different, explaining our inherent dependence on $K$ when $\sigma_{1}>0$. Specifically, they assume a reward function $r(x,a;\theta)=\varphi(x,a)^{\top}\theta$, with a shared $\theta\in\mathbb{R}^{d}$ and a known mapping $\varphi$ (i.e., the setting in Remark˜3.1). In contrast, we consider $r(x,a;\theta)=x^{\top}\theta_{a}$, where $\theta=(\theta_{a})_{a\in[K]}\in\mathbb{R}^{dK}$, requiring the learning of $K$ separate $d$-dimensional action parameters. Using our proof techniques, we can show that dTS’s regret is independent of $K$ in their setting, assuming access to $\varphi$ (Remark˜3.1 explains that our algorithm and analysis extend nicely to this setting).
@@ 6 Conclusion
We use a pre-trained diffusion model as a strong and flexible prior for dTS. Diffusion pre-training leverages abundant offline data, which is subsequently refined through online interactions via our tractable posterior approximation. This approximation enables efficient posterior sampling and updates while maintaining strong empirical performance. Moreover, dTS admits a simple Bayesian regret bound in the linear-Gaussian setting. Broader impact and computational considerations are discussed in Appendices˜I and J.
As outlined in Appendix˜H, extending our theoretical guarantees beyond the linear-Gaussian case, establishing robustness to model misspecification, and exploring $K$-independent or offline RL extensions constitute key future directions. Ultimately, this framework bridges online decision-making and generative modeling, offering a principled path toward diffusion model fine-tuning. Indeed, this work can be viewed as online diffusion model fine-tuning using implicit interaction data.
@@ Appendix A Extended related work
Large action spaces. Roughly speaking, the regret bound of dTS scales with $K\sigma_{1}^{2}$ rather than $K\sum_{\ell}\sigma_{\ell}^{2}$. This is particularly beneficial when $\sigma_{1}$ is small, a common scenario in diffusion models with decreasing variances. A notable case is when $\sigma_{1}=0$, where the regret becomes independent of $K$. Also, our analysis (Section˜4.1) indicates that the gap in performance between dTS and LinTS becomes more pronounced when the number of action increases, highlighting dTS’s suitability for large action spaces. Note that some prior works [25, 67, 70] proposed bandit algorithms that do not scale with $K$. However, our setting differs significantly from theirs, explaining our inherent dependency on $K$ when $\sigma_{1}>0$. Precisely, they assume a reward function of $r(x,a)=\phi(x,a)^{\top}\theta$, with a shared $\theta\in\mathbb{R}^{d}$ across actions and a known mapping $\phi$. In contrast, we consider $r(x,a)=x^{\top}\theta_{a}$, requiring the learning of $K$ separate $d$-dimensional action parameters. In their setting, with the availability of $\phi$, the regret of dTS would similarly be independent of $K$. However, obtaining such a mapping $\phi$ can be challenging as it needs to encapsulate complex context-action dependencies. Notably, our setting reflects a common practical scenario, such as in recommendation systems where each product is often represented by its embedding. In summary, the dependency on $K$ is more related to our setting than the method itself, and dTS would scale with $d$ only in their setting. Note that dTS is both computationally and statistically efficient (Section˜4.1). This becomes particularly notable in large action spaces. Our empirical results in Fig.˜1, notably with $K=10^{4}$, demonstrate that dTS significantly outperforms the baselines. More importantly, the performance gap between dTS and these baselines is larger when the number of actions ($K$) increases, highlighting the improved scalability of dTS to large action spaces.
@@ B.2 Posterior derivation in the linear diffusion case
$p(H_{t,a}\mid\theta_{a})\;\approx\;\mathcal{N}\!\big(\theta_{a};\hat{B}_{t,a},\hat{G}_{t,a}^{-1}\big),$
@@ Appendix C Posterior derivations for non-linear diffusion models
After deriving the exact posteriors for the linear link function case (Section˜B.2), we now return to the general setting with arbitrary, potentially non-linear link functions $f_{\ell}$. In this case, approximation is required since both the link functions and the reward distribution may be non-linear. To not lose the potentially high expressiveness of non-linearity, we do not approximate the overall posterior by a Gaussian. Instead, we approximate its components: namely, the local conditional densities $p(\theta_{a}\mid\psi_{1},H_{t,a})$ and $p(\psi_{\ell-1}\mid\psi_{\ell},H_{t})$, using the closed-form Gaussian expressions derived in Section˜B.2, while preserving the hierarchical diffusion structure. This yields an approximate posterior that remains a diffusion model but with updated, data-dependent means and covariances.
Concretely, two modifications are introduced. First, each linear term $\mathrm{W}_{\ell}\psi_{\ell}$ in Eq.˜13 is replaced by $f_{\ell}(\psi_{\ell})$, reflecting that the prior mean is now $f_{\ell}(\psi_{\ell})$ rather than $\mathrm{W}_{\ell}\psi_{\ell}$. This captures the distinction between the linear diffusion model in Eq.˜10 and the general, potentially non-linear, model in Eq.˜1. Second, matrix multiplications involving $\mathrm{W}_{\ell}$ in Eq.˜15 and Eq.˜16 are omitted.
Despite its simplicity, this approximation is computationally efficient and avoids the expensive posterior sampling required by more elaborate methods. Deriving the exact linear-case posteriors was therefore essential not only for theoretical analysis but also for enabling this practical and scalable extension. Moreover, the approximation preserves key properties of the exact posterior: it recovers the prior in the absence of data, and as data accumulates, the prior’s influence naturally diminishes. Finally, although Gaussian forms are used locally within each conditional factor, the global posterior remains non-Gaussian and retains the diffusion structure of the prior, but with updated means and covariances.
@@ D.1 Additional discussion: link to two-level hierarchies
The linear diffusion Eq.˜10 can be marginalized into a 2-level hierarchy using two different strategies. The first one yields,
@@
where $\Omega_{2}=\sum_{\ell=1}^{L}\sigma_{\ell+1}^{2}\mathrm{B}_{\ell}\mathrm{B}_{\ell}^{\top}$. Recently, HierTS [33] was developed for such two-level graphical models, and we call HierTS under Eq.˜17 by HierTS-1 and HierTS under Eq.˜18 by HierTS-2. Then, we start by highlighting the differences between these two variants of HierTS. First, their regret bounds scale as
@@
When $K\approx L$, the regret bounds of HierTS-1 and HierTS-2 are similar. However, when $K>L$, HierTS-2 outperforms HierTS-1. This is because HierTS-2 puts more uncertainty on a single $d$-dimensional latent parameter $\psi_{1}$, rather than $K$ individual $d$-dimensional action parameters $\theta_{a}$. More importantly, HierTS-1 implicitly assumes that action parameters $\theta_{a}$ are conditionally independent given $\psi_{L}$, which is not true. Consequently, HierTS-2 outperforms HierTS-1. Note that, under the linear diffusion model Eq.˜10, dTS and HierTS-2 have roughly similar regret bounds. Specifically, their regret bounds dependency on $K$ is identical, where both methods involve multiplying $K$ by $\sigma_{1}^{2}$, and both enjoy improved performance compared to HierTS-1. That said, note that Theorems˜E.1 and E.2 provide an understanding of how dTS’s regret scales under linear link functions $f_{\ell}$, and do not say that using dTS is better than using HierTS when the link functions $f_{\ell}$ are linear since the latter can be obtained by a proper marginalization of latent parameters (i.e., HierTS-2 instead of HierTS-1). While such a comparison is not the goal of this work, we still provide it for completeness next.
@@ F.1 Sketch of the proof
Applying Lemma˜F.1 requires proving that the marginal action-posterior densities of $\theta_{a}\mid H_{t}$ in Eq.˜2 are Gaussian and computing their covariances, while we only know the conditional action-posteriors $p(\theta_{a}\mid\psi_{1},H_{t})$ and latent-posteriors $p(\psi_{\ell-1}\mid\psi_{\ell},H_{t})$. This is achieved by leveraging the preservation properties of the family of Gaussian distributions [38] and the total covariance decomposition [66] which leads to the next lemma.
@@
The marginal covariance matrix $\check{\Sigma}_{t,a}$ in Eq.˜23 decomposes into $L+1$ terms. The first term corresponds to the posterior uncertainty of $\theta_{a}\mid\psi_{1}$. The remaining $L$ terms capture the posterior uncertainties of $\psi_{L}$ and $\psi_{\ell-1}\mid\psi_{\ell}$ for $\ell\in[L]/\{1\}$. These are then used to quantify the posterior information gain of latent parameters after one round as follows.
@@
Finally, Lemma˜F.2 is used to decompose $\|X_{t}\|_{\check{\Sigma}_{t,A_{t}}}^{2}$ in Eq.˜22 into $L+1$ terms. Each term is bounded thanks to Lemma˜F.3. This results in the Bayes regret bound in Theorem˜E.1.
@@ F.2 Technical contributions
In this proof, we heavily rely on the total covariance decomposition [66]. Also, refer to [33, Section 5.2] for a brief introduction to this decomposition. Now, from Eq.˜12, we have that
@@
However, $\bar{\bar{\Sigma}}_{t,1}=\mathrm{cov}\left[\psi_{1}\,\middle|\,H_{t}\right]$ is different from $\bar{\Sigma}_{t,1}=\mathrm{cov}\left[\psi_{1}\,\middle|\,H_{t},\psi_{2}\right]$ that we already derived in Eq.˜13. Thus we do not know the expression of $\bar{\bar{\Sigma}}_{t,1}$. But we can use the same total covariance decomposition trick to find it. Precisely, let $\bar{\bar{\Sigma}}_{t,\ell}=\mathrm{cov}\left[\psi_{\ell}\,\middle|\,H_{t}\right]$ for any $\ell\in[L]$. Then we have that
@@
(I) Base case. Let $u=\sigma^{-1}\hat{\Sigma}_{t,A_{t}}^{\frac{1}{2}}X_{t}$ From the expression of $\bar{\Sigma}_{t,1}$ in Eq.˜13, we have that
@@
where we use that by definition of $\sigma_{\textsc{max}}^{2}$ in Lemma˜F.3, we have that $\sigma_{\textsc{max}}^{2}\geq 1+\sigma^{-2}\sigma_{1}^{2}$. Therefore, by taking the inverse, we get that $\frac{1}{1+u^{\top}u}\geq\sigma_{\textsc{max}}^{-2}$. Combining this with Eq.˜26 leads to
@@
where $(i)$ follows from Eq.˜28, and we use the following inequality in $(ii)$
@@
Therefore, from Section˜F.5 and Eq.˜27, we get that
@@
The last inequality follows from the expression of $\bar{\Sigma}_{n+1,\ell}^{-1}$ in Eq.˜13 that leads to
@@ G.4 CIFAR-10 ablation study
CIFAR-10. In Fig.˜3(a) in Section˜5.2, we showed that with only 10 pre-training samples, dTS outperforms LinTS on the Swiss-roll benchmark. We now extend this analysis to the vision dataset CIFAR-10 [40] (similar results were obtained on MNIST [48]). Our setting is similar to that in Hong et al. [32] and we use dTS’s variant that uses a single shared parameter $\theta\in^{d}$ (Remark˜3.1) because it is more suited for this setting. These additional ablations on CIFAR-10 confirm that dTS consistently benefits from offline pre-training, even when the true prior is not a diffusion model. Specifically, we vary the percentage of offline data used to train the prior and compare against both HierTS and LinTS.
@@ Appendix H Extensions
$K$-independent regret. In the setting of Remark˜3.1, where $r(x,a;\theta)=\varphi(x,a)^{\top}\theta$ and $\theta$ is shared across actions, our proof techniques imply $K$-independent regret once $\varphi$ is known or accurately estimated. This connects to structured large-action-space results where regret does not scale with $K$ [25, 67, 70]. Exploring further the use of diffusion models in such setting is promising.

### v3

@@ 4.1 Discussion
Regret independent of $K$? dTS’s regret bound scales with $K\sigma_{1}^{2}$ rather than $K\sum_{\ell}\sigma_{\ell}^{2}$, which is particularly advantageous when $\sigma_{1}$ is small, as is often the case with diffusion model priors. Both our theoretical bound and empirical results show that dTS’s advantage over LinTS increases as the action space grows. Nevertheless, dTS’s regret still depends on $K$; this dependence arises from the problem setting rather than from the algorithm itself. Prior works [25, 67, 70] have proposed bandit algorithms whose regret does not scale with $K$. This difference stems from the setting considered: we study the disjoint (per-action) case $r(x,a;\theta)=x^{\top}\theta_{a}$, where $\theta=(\theta_{a})_{a\in[K]}\in\mathbb{R}^{dK}$, requiring the learning of $Kd$ parameters and thus introducing an inherent dependence on $K$ when $\sigma_{1}>0$. In contrast, $K$-independent regret results are obtained in the shared-parameter setting described in Remark˜2.1, where $r(x,a;\theta)=\varphi(x,a)^{\top}\theta$ and only a single $d$-dimensional parameter must be learned. However, this formulation requires access to the feature map $\varphi$. Fortunately, dTS is also compatible with this setting (Section˜3.3), in which case its regret would indeed be independent of $K$.
@@ 6 Conclusion
We use a pre-trained diffusion model as a strong and flexible prior for dTS. Diffusion pre-training leverages abundant offline data, which is then fine-tuned through online interactions via our tractable posterior approximation. This approximation enables efficient posterior sampling and updates while maintaining strong empirical performance. Moreover, dTS admits a simple Bayesian regret bound in the linear–Gaussian setting. Broader impact and computational considerations are discussed in Appendices˜I and J, and directions for future work are provided in Appendix˜H.
@@ Appendix A Extended related work
Large action spaces. The regret bound of dTS scales with $K\sigma_{1}^{2}$ rather than $K\sum_{\ell}\sigma_{\ell}^{2}$, which is particularly advantageous when $\sigma_{1}$ is small: a common case in diffusion models with decreasing layer variances. In the limiting case $\sigma_{1}=0$, the regret becomes independent of $K$. Our theoretical analysis (Section˜4.1) and empirical results both show that the performance gap between dTS and LinTS widens as the number of actions increases, highlighting dTS’s suitability for large action spaces.
Some prior works [25, 67, 70] achieve regret bounds that do not scale with $K$. This discrepancy stems from the problem setting rather than the algorithm itself. Specifically, those works adopt a shared-parameter model $r(x,a)=\varphi(x,a)^{\top}\theta$ with a single parameter $\theta\in\mathbb{R}^{d}$ and a known feature map $\varphi$, whereas we study the disjoint case $r(x,a)=x^{\top}\theta_{a}$ with $K$ separate $d$-dimensional parameters. In the shared-parameter setting (see Remarks˜2.1 and 3.3), dTS would similarly achieve regret independent of $K$.
In summary, the dependence on $K$ arises from the modeling choice rather than a limitation of dTS. When $\varphi$ is available, dTS scales only with $d$; otherwise, in the per-action setting, it remains both computationally and statistically efficient (Section˜4.1). Empirically, even for very large action spaces (e.g., $K=10^{4}$), dTS substantially outperforms existing baselines, with the performance gap increasing as $K$ grows: highlighting its scalability to large action spaces.
@@ B.2 Posterior derivation in the linear diffusion case
$p(H_{t,a}\mid\theta_{a})\;\approx\;\mathcal{N}\!\big(\theta_{a};\hat{B}_{t,a},\hat{G}_{t,a}^{-1}\big),$ (19)
@@ Appendix C Posterior derivations for non-linear diffusion models
The general diffusion model (Eq.˜1, which is our case of interest) involves two sources of non-linearity: (i) the reward distribution $P(\cdot\mid x;\theta)$, which may follow a non-linear generalized linear model (GLM), and (ii) the diffusion links $f_{\ell}(\psi_{\ell})$, which can be arbitrary non-linear functions. Both sources make the posterior intractable, and therefore two approximations are needed.
First approximation (likelihood). We first approximate the reward likelihood by a Gaussian density (as we did above in Eq.˜19). After this substitution, the model becomes conditionally Gaussian given the latent variables. This step is exact when the reward model is linear-Gaussian, and approximate otherwise.
Second approximation (diffusion hierarchy). Even after the likelihood is approximated, the diffusion hierarchy remains non-linear because of the non-linear mappings $f_{\ell}$. To handle this, we reuse the exact Gaussian posteriors derived for the linear diffusion case (Section˜B.2) and generalize them as follows:
  • Replace each linear mapping $\mathrm{W}_{\ell}\psi_{\ell}$ by its non-linear counterpart $f_{\ell}(\psi_{\ell})$, which represents the mean of the diffusion prior at layer $\ell$.
  • Remove matrix multiplications involving $\mathrm{W}_{\ell}$ in the recursive updates.
This step can be viewed as extending the linear-Gaussian posterior formulas to a general non-linear setting, without performing explicit linearization or optimization.
Resulting approximation. The two steps above yield a posterior where each conditional factor $p(\theta_{a}\mid\psi_{1},H_{t,a})$ and $p(\psi_{\ell-1}\mid\psi_{\ell},H_{t})$ remains Gaussian with updated means and covariances, while the overall model retains the hierarchical diffusion structure. The approximation satisfies two desirable properties: it exactly recovers the diffusion prior when no data is available, and as more data is observed, the likelihood terms dominate and the prior influence fades naturally.
This approach is computationally efficient, avoids costly posterior sampling or variational optimization, and remains expressive since the overall posterior is still a diffusion model with non-linear link functions and covariances that are now data-dependent.
@@ D.1 Additional discussion: link to two-level hierarchies
The linear diffusion Eq.˜17 can be marginalized into a 2-level hierarchy using two different strategies. The first one yields,
@@
where $\Omega_{2}=\sum_{\ell=1}^{L}\sigma_{\ell+1}^{2}\mathrm{B}_{\ell}\mathrm{B}_{\ell}^{\top}$. Recently, HierTS [33] was developed for such two-level graphical models, and we call HierTS under Eq.˜25 by HierTS-1 and HierTS under Eq.˜26 by HierTS-2. Then, we start by highlighting the differences between these two variants of HierTS. First, their regret bounds scale as
@@
When $K\approx L$, the regret bounds of HierTS-1 and HierTS-2 are similar. However, when $K>L$, HierTS-2 outperforms HierTS-1. This is because HierTS-2 puts more uncertainty on a single $d$-dimensional latent parameter $\psi_{1}$, rather than $K$ individual $d$-dimensional action parameters $\theta_{a}$. More importantly, HierTS-1 implicitly assumes that action parameters $\theta_{a}$ are conditionally independent given $\psi_{L}$, which is not true. Consequently, HierTS-2 outperforms HierTS-1. Note that, under the linear diffusion model Eq.˜17, dTS and HierTS-2 have roughly similar regret bounds. Specifically, their regret bounds dependency on $K$ is identical, where both methods involve multiplying $K$ by $\sigma_{1}^{2}$, and both enjoy improved performance compared to HierTS-1. That said, note that Theorems˜E.1 and E.2 provide an understanding of how dTS’s regret scales under linear link functions $f_{\ell}$, and do not say that using dTS is better than using HierTS when the link functions $f_{\ell}$ are linear since the latter can be obtained by a proper marginalization of latent parameters (i.e., HierTS-2 instead of HierTS-1). While such a comparison is not the goal of this work, we still provide it for completeness next.
@@ F.1 Sketch of the proof
Applying Lemma˜F.1 requires proving that the marginal action-posterior densities of $\theta_{a}\mid H_{t}$ in Eq.˜3 are Gaussian and computing their covariances, while we only know the conditional action-posteriors $p(\theta_{a}\mid\psi_{1},H_{t})$ and latent-posteriors $p(\psi_{\ell-1}\mid\psi_{\ell},H_{t})$. This is achieved by leveraging the preservation properties of the family of Gaussian distributions [38] and the total covariance decomposition [66] which leads to the next lemma.
@@
The marginal covariance matrix $\check{\Sigma}_{t,a}$ in Eq.˜31 decomposes into $L+1$ terms. The first term corresponds to the posterior uncertainty of $\theta_{a}\mid\psi_{1}$. The remaining $L$ terms capture the posterior uncertainties of $\psi_{L}$ and $\psi_{\ell-1}\mid\psi_{\ell}$ for $\ell\in[L]/\{1\}$. These are then used to quantify the posterior information gain of latent parameters after one round as follows.
@@
Finally, Lemma˜F.2 is used to decompose $\|X_{t}\|_{\check{\Sigma}_{t,A_{t}}}^{2}$ in Eq.˜30 into $L+1$ terms. Each term is bounded thanks to Lemma˜F.3. This results in the Bayes regret bound in Theorem˜E.1.
@@ F.2 Technical contributions
In this proof, we heavily rely on the total covariance decomposition [66]. Also, refer to [33, Section 5.2] for a brief introduction to this decomposition. Now, from Eq.˜20, we have that
@@
However, $\bar{\bar{\Sigma}}_{t,1}=\mathrm{cov}\left[\psi_{1}\,\middle|\,H_{t}\right]$ is different from $\bar{\Sigma}_{t,1}=\mathrm{cov}\left[\psi_{1}\,\middle|\,H_{t},\psi_{2}\right]$ that we already derived in Eq.˜21. Thus we do not know the expression of $\bar{\bar{\Sigma}}_{t,1}$. But we can use the same total covariance decomposition trick to find it. Precisely, let $\bar{\bar{\Sigma}}_{t,\ell}=\mathrm{cov}\left[\psi_{\ell}\,\middle|\,H_{t}\right]$ for any $\ell\in[L]$. Then we have that
@@
(I) Base case. Let $u=\sigma^{-1}\hat{\Sigma}_{t,A_{t}}^{\frac{1}{2}}X_{t}$ From the expression of $\bar{\Sigma}_{t,1}$ in Eq.˜21, we have that
@@
where we use that by definition of $\sigma_{\textsc{max}}^{2}$ in Lemma˜F.3, we have that $\sigma_{\textsc{max}}^{2}\geq 1+\sigma^{-2}\sigma_{1}^{2}$. Therefore, by taking the inverse, we get that $\frac{1}{1+u^{\top}u}\geq\sigma_{\textsc{max}}^{-2}$. Combining this with Eq.˜34 leads to
@@
where $(i)$ follows from Eq.˜36, and we use the following inequality in $(ii)$
@@
Therefore, from Section˜F.5 and Eq.˜35, we get that
@@
The last inequality follows from the expression of $\bar{\Sigma}_{n+1,\ell}^{-1}$ in Eq.˜21 that leads to
@@ G.4 CIFAR-10 ablation study
CIFAR-10. In Fig.˜3(a) in Section˜5.2, we showed that with only 10 pre-training samples, dTS outperforms LinTS on the Swiss-roll benchmark. We now extend this analysis to the vision dataset CIFAR-10 [40] (similar results were obtained on MNIST [48]). Our setting is similar to that in Hong et al. [32] and we use dTS’s variant that uses a single shared parameter $\theta\in^{d}$ (Remarks˜2.1 and 3.3) because it is more suited for this setting. These additional ablations on CIFAR-10 confirm that dTS consistently benefits from offline pre-training, even when the true prior is not a diffusion model. Specifically, we vary the percentage of offline data used to train the prior and compare against both HierTS and LinTS.
@@ Appendix H Extensions
$K$-independent regret. In the setting of Remark˜2.1, where $r(x,a;\theta)=\varphi(x,a)^{\top}\theta$ and $\theta$ is shared across actions, our proof techniques imply $K$-independent regret once $\varphi$ is known or accurately estimated. This connects to structured large-action-space results where regret does not scale with $K$ [25, 67, 70]. Exploring further the use of diffusion models in such setting is promising.